# SimuRA: Towards General Goal-Oriented Agent via Simulative Reasoning Architecture with LLM-Based World Model

### v1

@@ 5 Limitations
Due to the modular pipeline and thorough exploration of multiple plans in world model planning, the current agent takes longer than typical LLM agents to run. Speeding up world-model-based reasoning with appropriate caching and parallelization strategies is an important part of our future work. Agent capabilities can be limited by the tooling. For example, with open-source browser environments, web agents are often blocked by Captcha or anti-scraping tools from certain websites. Deeper integration with user browser can help solve this issue. As agent-based automation become more integrated into browsing and computer-use workflows, we also encourage conversations around fair use and protocols around agent access of certain websites. We are currently only using the text portion of the webpage observations, which can miss information like images and layout information (e.g., occlusions). While existing work are experimenting with visual-based web browsing, it is still challenging to combine multimodal perception and planning, which we are excited to keep working on.
@@ 6 Conclusion
In this paper, we have presented SimuRA, a general goal-oriented architecture for optimal agent decision-making. Empowered by simulation-based planning using world model and modeling of agent internal belief space activities using natural language as latent representation, we see significant and strong improvements on a range of tasks in web browsing experiments, with world model-based reasoning showing improved reasoning capacity compared to LLM autoregressive reasoning.
We are very excited about the possibilities for a single, general, superintelligent agent, but are also keenly aware of the risks for individuals and societies. On the capability side, we aim to test on more types of environments (e.g., software development) and continue developing functional components that strengthen the agent (e.g., multi-agent interaction and long-term memory). On the safety / alignment side, we look forward to engaging the community in discussions about how to ensure such an agent stays aligned with our shared values, priorities, and welfare.
Acknowledgement
This work was supported in part by the Samsung GRO Project “Efficient Designs for Generative and Agent LLM Development.” We thank Zhoujun Cheng, Shibo Hao, and Xinyu Pi from MixLab; Han Guo, Nicholas Ho, and Bowen Tan from SAILING Lab; Li Erran Li from AWS, Zora Wang from NeuLab; and Sarah Cheah and Hector Ren from MBZUAI for their insightful feedback and discussions. We are also grateful for their helpful suggestions throughout the project. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of Samsung.
@@ Appendix B Prompts for Web Browsing Implementation
Describe all the elements in the current webpage observation. Note any dialogs, progress indicators, or error messages. Include any interactive elements and their values or if they are blank. Note any detailed information such as facts, entities, or data that are relevant to the task. Report any error messages like whether the last action was correct. Try to be as comprehensive and detailed as possible.
@@
Describe all the elements in the webpage after the agent attempts to carry out the intent. Note that the execution may not be successful, so you will have to infer the result of the action. Note any dialogs, progress indicators, or error messages. Include any interactive elements and their values or if they are blank. Note any detailed information such as facts, entities, or data that are relevant to the task. Report any error messages displayed. Try to be as comprehensive and detailed as possible.

### v2

@@ 5 Limitations
Due to the modular pipeline and thorough exploration of multiple plans in world-model-based simulation, the current agent takes longer than typical reactive agents to run. Speeding up world-model-based reasoning with appropriate caching and parallelization strategies is an important part of our future work. In addition, agent capabilities can be limited by the tooling. For example, with open-source browser environments, web agents are often blocked by Captcha or anti-scraping tools from certain websites. Deeper integration with user browser can help solve this issue. As agent-based automation become more integrated into browsing and computer-use workflows, we also encourage conversations around fair use and protocols around agent access of certain websites. We are currently only using the text portion of the webpage observations, which can miss information like images and layout information (e.g., occlusions). While existing work are experimenting with visual-based web browsing, it is still challenging to combine multimodal perception and planning, which we are excited to keep working on.
@@ 6 Conclusion
In this paper, we have presented SimuRA, a goal-oriented architecture for agentic reasoning and decision-making. By augmenting autoregressive reasoning with simulation-based planning through a world model and by representing internal belief states in the semantically structured latent space of natural language, SimuRA demonstrates consistent improvements across a range of web-interaction tasks. In particular, our results suggest that explicit world-model reasoning can enhance the planning and reasoning capacity of agents beyond what purely autoregressive reasoning provides, underscoring the value of simulative reasoning as a core component of general intelligence.
Looking ahead, we see SimuRA as a foundation for developing more capable, robust, and controllable agentic systems. On the capability side, we aim to test on more types of environments (e.g., embodied sandboxes and physical space) and to continue developing functional components that including multi-agent interaction and long-term memory. On the safety and alignment side, we look forward to engaging the community in discussions about how explicit world-model reasoning may help build agents that remain transparent and aligned with human values and collective welfare.
Acknowledgment
This work was supported in part by the Samsung GRO Project “Efficient Designs for Generative and Agent LLM Development.” We thank Graham Neubig and Zora Wang from NeuLab; Yilin Shen and Hongxia Jin from Samsung Research; Zhoujun Cheng, Shibo Hao, and Xinyu Pi from MixLab; Han Guo, Nicholas Ho, and Bowen Tan from SAILING Lab; Li Erran Li from AWS, and Sarah Cheah and Hector Ren from MBZUAI for their insightful feedback and discussions. We are also grateful for their helpful suggestions throughout the project. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of Samsung.
@@ Appendix B Prompts for Web Browsing Implementation
Describe all the elements in the current webpage observation. Note any dialog, progress indicators, or error messages. Include any interactive elements and their values or if they are blank. Note any detailed information such as facts, entities, or data that are relevant to the task. Report any error messages like whether the last action was correct. Try to be as comprehensive and detailed as possible.
@@
Describe all the elements in the webpage after the agent attempts to carry out the intent. Note that the execution may not be successful, so you will have to infer the result of the action. Note any dialog, progress indicators, or error messages. Include any interactive elements and their values or if they are blank. Note any detailed information such as facts, entities, or data that are relevant to the task. Report any error messages displayed. Try to be as comprehensive and detailed as possible.